A body of mass 1 kg is thrown vertically upward and the body comes back on the earth surface after 8 s. Find the K. E. and potential energy after 2 s. g = 10 ms².


1
Expert's answer
2021-11-01T19:24:37-0400

"m = 1kg"

"t_1 = 8s"

"t_2 = 2s"

"g =10\frac{m}{s^2}"

"h(t) = v_0t - \frac{gt^2}{2}"

"h(t_1)=0"

"8v_0-\frac{10*8^2}{2}=0"

"v_0 = 40\frac{m}{s}"

"v(t) = v_0-gt"

"v(t_2) = 40-10*2=20\frac{m}{s}"

"K.E= \frac{mv^2}{2}=\frac{1*20^2}{2}=200J"

"h(t_2)= 40*2-\frac{10*2^2}{2}=60m"

"P.E = mgh = 1*10*60= 600J"


"\text{Answer: }K.E=200J;P.E=600J"